OpenStudy (anonymous):

h(x)+ x to the second power - 5.. find h(-3) -14 15 -11 4

OpenStudy (zzr0ck3r):

? can you rewrite it just as the book has it

OpenStudy (anonymous):

Can you write the question once again in more understandable form so that we can understand it and help you..

OpenStudy (anonymous):

if h(x) = x^2 - then find h(-3) thats exactly what the question says i don't understand it at all/

OpenStudy (anonymous):

-5*

jimthompson5910 (jim_thompson5910):

h(x) = x^2 - 5 Replace x with -3 to get h(-3) = (-3)^2 - 5 Now square -3 to get (-3)^2 = (-3)*(-3) = ???

OpenStudy (anonymous):

9?

jimthompson5910 (jim_thompson5910):

Good

jimthompson5910 (jim_thompson5910):

So h(-3) = (-3)^2 - 5 becomes h(-3) = 9 - 5

OpenStudy (anonymous):

would it be 4?!

jimthompson5910 (jim_thompson5910):

You nailed it

OpenStudy (anonymous):

your are a genius.

jimthompson5910 (jim_thompson5910):

You did most the work lol

jimthompson5910 (jim_thompson5910):

but thanks

OpenStudy (anonymous):

Your welcome!!(: